OpenStudy (anonymous):

find all solutions: 25 - 16cube root of f + 1 = -7

OpenStudy (anonymous):

Isolate the cube root term by subtracting 25 from both sides and dividing both sides by -16 \[\sqrt[3]{f+1}=2\]

OpenStudy (anonymous):

Then, cube both sides. f+1 = 8 and solve for f f = 7

OpenStudy (anonymous):

okay thanks alot...my next question is 4x^2 cube root of 3 times 2x cube root of 3